Estimate the order of magnitude of the velocity of sound in the metal.


1
Expert's answer
2021-09-29T09:54:51-0400

The speed of sound can be expressed in terms of density (order of magnitude 1000) and young's modulus of metal (order of magnitude 109:


"OM_v=\sqrt\frac{OM_E}{OM_\rho}=\sqrt{\frac{10^9}{1000}}=10^3."